OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1=13 and a17=-51

OpenStudy (anonymous):

Arithmetic sequences have constant differecne so a17-a1=-64 17 - 1 = 16, so you know how many terms are there between the two given, so the difference in two consecutive terms is -64/16=-4 If you want to know to a35 from a17 you know there are, 35-17=18, terms, just -4*18=-72 Then a35=a17-72=-51-72=-123

OpenStudy (anonymous):

Thank you that is a great help and your explaination really helped me to understand

OpenStudy (anonymous):

Yw (: